Depreciation for Partial Periods Norman Delivery Company purchased a new delivery truck for $66,600 on April 1, 2016. The truck is expected to have a service life of 5 years or 108,000 miles and a residual value of $5,520. The truck was driven 9,300 miles in 2016 and 11,200 miles in 2017. Norman computes depreciation to the nearest whole month.

Required: Compute depreciation expense for 2016 and 2017 using the For interim computations, carry amounts out to two decimal places. Round your final answer to the nearest dollar.

Straight-line method

2016 $

2017 $

Sum-of-the-years'-digits method

2016

2017

Double-declining-balance method

2016

2017

Activity method

2016

2017

For each method, what is the book value of the machine at the end of 2016? At the end of 2017? (Round your answers to the nearest dollar.)

Straight-line method

2016

2017

Sum-of-the-years'-digits method

2016 2017

Double-declining-balance method

2016

2017

Activity method

2016

2017
Answer #1
Depreciable cost = Cost - Residual value = $66600 - $5520 = $61080
Depreciation expense:
Straight-line method
2016: $61080 x 1/5 x 9/12 = $9162
2017: $61080 x 1/5 = $12216
Sum-of-the-years'-digits method
Sum-of-the-years'-digits = 1 + 2 + 3 + 4 + 5 = 15
2016: $61080 x 5/15 x 9/12 = $15270
2017: ($61080 x 5/15 x 3/12) + ($61080 x 4/15 x 9/12) = $5090 + $12216 = $17306
Double-declining-balance method
Straight-line depreciation rate = 100%/5 years = 20%
Double-declining-balance depreciation rate = 2 x 20% = 40%
2016: 40% x $66600 x 9/12 = $19980
2017: 40% x ($66600 - $19980) = 40% x $46620 = $18648
Activity method
Depreciation per mile = $61080/108000 miles = $0.57 per mile
2016: 9300 x $0.57 = $5301
2017: 11200 x $0.57 = $6384
Book value:
Straight-line method
2016: $66600 - $9162 = $57438
2017: $57438 - $12216 = $45222
Sum-of-the-years'-digits method
2016: $66600 - $15270 = $51330
2017: $51330 - $17306 = $34024
Double-declining-balance method
2016: $66600 - $19980 = $46620
2017: $46620 - 18648 = $27972
Activity method
2016: $66600 - $5301 = $61299
2017: $61299 - $6384 = $54915